OpenStudy (otherworldly):

Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%. Roger has 40 cars in his collection. Each year he collects 1 additional car. Part A: Write functions to represent Cory and Roger's collections throughout the years. (4 points)

OpenStudy (otherworldly):

Would the function for cory be f(x) = 15(1.2)^x and Roger be g(x) = 40 + 1x

OpenStudy (otherworldly):

@steve816 can you tell me if i am correct

OpenStudy (otherworldly):

@jhonyy9

OpenStudy (otherworldly):

@3mar do u think i did something incorrect

OpenStudy (3mar):

Give me 3 min, please!

OpenStudy (3mar):

Sorry for being late! "Cory has 15 die-cast cars in his collection. Each year his collection increases by 20%." That means Cory has an \(\color{red}{initial~ amount}\) of \(\color{red}{15}\) cars and the rate of growing\increasing is \(\color{red}{0.2}\),,,so that the relation would be an exponential growing function as follows: